### v2

@@ Abstract
Knowledge graphs have been widely adopted in both enterprises, such as the Google Knowledge Graph, and open platforms like Wikidata, to represent domain knowledge and support artificial intelligence applications. They model real-world information as nodes and edges. To embrace flexibility, knowledge graphs often lack enforced schemas (i.e., ontologies), leading to potential data quality issues, such as semantically overlapping nodes. Yet ensuring their quality is essential, as issues in the data can affect applications relying on them. To assess the quality of knowledge graphs, existing works propose either high-level frameworks comprising various data quality dimensions without concrete implementations, define tools that measure data quality with ad-hoc SPARQL queries, or promote the usage of constraint languages, such as the Shapes Constraint Language (SHACL), to assess and improve the quality of the graph. Although the latter approaches claim to address data quality assessment, none of them comprehensively tries to cover all data quality dimensions. In this paper, we explore this gap by investigating the extent to which SHACL can be used to assess data quality in knowledge graphs. Specifically, we defined SHACL shapes for 69 data quality metrics proposed by Zaveri et al. [1] and implemented a prototype that automatically instantiates these shapes and computes the corresponding data quality measures from their validation results. All resources are provided for repeatability.
@@
Knowledge graphs (KGs) have been increasingly used to represent domain knowledge and support artificial intelligence applications, such as information retrieval and question answering [2]. As a result, ensuring the quality of KGs is crucial for applications that rely on their input [3]. Major technology companies, such as Google (Google Knowledge Graph [4]) and Amazon (Alexa Knowledge Graph [5]), have developed proprietary KGs [2], while collaborative KGs, such as Wikidata [6], DBpedia [7] and YAGO [8], have emerged as open-source alternatives. This growing adoption of KGs is reflected in the growth of the Linked Open Data (LOD) cloud, which has enabled the publication and interlinking of numerous datasets across domains, with 1 656 resources as of Nov. 2024 [9].
@@
In recent years, constraint languages, such as the Shapes Constraint Language (SHACL) [25] and Shapes Expression Language (ShEx) [26], have emerged to enable RDF graph validation by specifying constraints as shapes [10]. Despite syntactic differences, both languages enable the definition of constraints on nodes and their value nodes (i.e., values reachable via properties or paths), allowing for the detection of data violations [27]. These languages enable a low-level perspective on DQ, focusing on concrete error detection through constraint checks. But unlike ad-hoc SPARQL-based approaches, constraint languages like SHACL offer a formal way for expressing validation rules.
@@
This paper explores how SHACL can be leveraged for DQA. In particular, we want to investigate the extent to which we can connect the high-level view on DQ dimensions with the low-level view on DQ, which entails identifying constraint violations using SHACL shapes. Thus, the contributions of this paper are as follows:
@@ 2 Related Work
In addition, several tools have been proposed over the years to assess the quality of KGs and linked data. Some of these tools focus on specific DQ dimensions [37, 38], while others try to cover a wider range of them [21, 23, 39, 40, 24, 41, 42, 43]. Moreover, some tools focus on assessing the quality of SPARQL endpoints [41, 38, 37], while others focus on the quality of the data itself [21, 23, 39, 40, 24, 42, 43]. Assessment is primarily done via ad-hoc SPARQL queries [23, 39, 37, 24, 41, 22, 43], while some introduce more complex techniques [21, 42]. Only [24] considers the usage of a constraint language. Additionally, some of these tools haven’t been maintained for some time or aren’t available. In Appendix A, we present a table summarizing these tools.
@@ 3 SHACL shapes for DQ dimensions
This paper builds on the comprehensive DQ metrics survey by Zaveri et al. [1], which groups dimensions into four categories. We follow their structure, defining SHACL core shapes for each metric or explaining SHACL core’s limitations when shapes cannot be defined. 111A discussion on SHACL extensions and the justification for the focus on SHACL core can be found in Section 5.3. We present the results of this study in tables 1 - 4, which illustrate the feasibility of implementing metrics from [1] with SHACL core. Symbols , p, and x denote the degree to which metric implementation was possible (full, partial, or not at all, respectively). For the shape definition, we made some realistic assumptions (A1–3):
@@
The ontology used is sufficiently defined: each class c and property p has triples c,rdf:type,rdfs:Class and p,rdf:type,rdf:Property, respectively. Each property p, includes domain and range triples p,rdfs:domain,d and p,rdfs:range,r. Other property characteristics (e.g., irreflexivity, asymmetric) are also defined. For every instance i of a class c defined in the ontology, i,rdf:type,owl:NamedIndividual exists.
@@
In tables 1 - 4, the use of assumptions is noted with , where X corresponds to the assumption number. In the following subsections, we summarize the coverage of SHACL core per group and discuss one representative DQ metric for a single dimension of each group. Whenever applicable, we provide the corresponding shape for the metric in Turtle syntax. The complete list of defined shapes is available in Appendix B.For each defined shape, we describe the validation result, and for the implemented ones, we indicate the DQ measure type: binary measure (1 if no violations, 0 otherwise), ratio measure (violation-based ratios), or composite measure (aggregated score across shape instances (properties or classes)).
@@ 3.1 Accessibility
Performance refers to how efficiently a system that hosts a large dataset can process data. For this dimension, Zaveri et al. [1] defined four metrics. We showcase the assessment of Performance with metric P1 in SHLABEL:shape:performance_hash_uris_entities. P1 was defined by [44] and checks for slash-URIs in datasets with over 500 000 triples. According to W3C recommendations, slash-URIs are preferred to identify entities in large graphs because hash (“#”) URIs can cause performance issues [45]. We follow the W3C Best Practices [46] and apply this rule to entities’ URIs. Here, SHLABEL:shape:performance_hash_uris_entities targets subjects using rdf:type with a regex pattern that does not allow # to appear in URIs. Note that the regex pattern identifies any hash occurrence (not just at the end of the URI) to cover cases such as https://www.example.org#Example1/, which still loads all entities with https://www.example.org# as base namespace. Therefore, the validation result will output nodes whose URI contains a hash in any part of the URI, not just at the end.
@@ 4 Prototype
This section presents our prototype for DQA using SHACL core. It is built with the Python libraries rdflib and PySHACL and provides a Streamlit dashboard to visualize results. The code is available on GitHub at https://github.com/caroocortes/SHACL-DQA-prototype.
Overview of the DQ assessment process. Our DQA process takes as input the data graph to be evaluated, optionally a metadata file (either the VoID or DCAT description of the graph), and an ontology, along with a set of vocabularies used in the data graph. Without an ontology, vocabularies, or a metadata file, fewer shapes will be instantiated, as some rely on these artifacts. Additionally, certain shapes will not be validated, such as those targeting void:Dataset (validated against the metadata file) or those checking for labels in classes and properties (validated against vocabularies/ontology). A configuration file allows users to customize preferred properties needed for SHACL shapes (e.g., we use rdf:type, rdfs:label, rdfs:comment, and owl:sameAs as default properties for defining shapes, but users may change this)
@@
After obtaining this information, we (2) instantiate the shapes’ templates and generate the shapes graph. Our prototype stores SHACL shapes as reusable templates with variables that are replaced with actual values during instantiation – a process where the template is populated with specific properties from either the graph profiling results or the configuration file of the dataset. For example, shape SHLABEL:shape:performance_hash_uris_entities uses the type property specified in the configuration file to create a concrete shape from its template form. Moreover, shape SHLABEL:shape:consistency_uniqueness_inverse_functional_property is instantiated with inverse-functional properties used in the dataset, obtained from the graph profiling results. Before shape validation, we perform a pre-processing step (based on assumption (A2)) to enrich the data graph with necessary triples. Specifically, we add type declarations for OWL-defined properties and classes defined in ontology/vocabularies, marking each property as an rdf:Property and each class as an rdfs:Class. We also explicitly type instances in vocabularies/ontology as owl:NamedIndividual to distinguish data entities from schema elements. For terms defined as rdfs:Datatype, we type them as rdfs:Class, in line with the RDFS entailment rule rdfs11. Additionally, when multiple vocabularies are reused, we include all relevant axioms in the data graph, as stated by the SHACL specification [25]. Then, we use the validator (3) provided by the PySHACL library to validate the shapes against the data graph, ontology, vocabularies, or metadata file, depending on the shape. Once the validation is completed, we (4) calculate the DQ measures from the validation result, which are later stored in a CSV file. Finally, the results can be visualized in a dashboard (5).
@@
Shape instantiation details. For the 69 metrics defined in Zaveri et al. [1], we defined 64 shapes. The mapping is not one-to-one: some metrics have no shape, others (e.g., V2 in Section 3.4) have multiple. We identified 38 shapes that could be instantiated without domain expert input, and excluded five more: one requiring merged graphs (SH11) and 4 using non-standard vocabularies (S13, S14, S53, S55) (see Appendix B for these shapes definitions). Of the 38, 11 rely on vocabulary or ontology terms, and are instantiated only with those found in the data graph (e.g., SHLABEL:shape:consistency_uniqueness_inverse_functional_property uses only inverse-functional properties present in the data). This avoids generating unnecessary shapes, as many data graphs may partially reuse vocabularies. The exceptions are the shapes for metrics CN2 and CP1: CN2 checks for property/class misuse by instantiating with all available classes and properties since misused ones do not appear in profiling results, while CP1 uses all defined classes in the vocabulary to check if they are used. See Appendix C for additional aspects of shape instantiation aimed at improving runtime efficiency.
@@
Evaluation. We evaluated the prototype with three datasets from the LOD Cloud (https://lod-cloud.net/): Temples of the Classical World (15 326 triples and 1 363 entities), DBTunes - John Peel Sessions (271 369 triples and 76 056 entities), and Drugbank (3 646 181 triples and 316 555 entities). Validation time ranged from 40 seconds to 3 hours, with 500 to 740 shapes instantiated depending on the dataset size. As we relied on an external library for the validation, we did not perform any runtime experiments. Table 5 presents the assessment results for the dataset Temples of the Classical World in detail. The results for the other datasets are provided on Github222https://github.com/caroocortes/SHACL-DQA-prototype. While Temples of the Classical World demonstrates strengths in aspects such as URI design and labeling, it exhibits several quality issues, particularly in metadata aspects, and correct vocabulary usage.
@@ 5 Suitability of SHACL core to assess RDF data quality
We now discuss the suitability of SHACL core to assess RDF DQ. Section 5.1 highlights where SHACL core falls short in covering certain DQ dimensions – metrics not mentioned are considered covered (see Section 3 and Appendix B). We then examine SHACL core’s limitations (Section 5.2), discuss SHACL core’s extensions (Section 5.3), and its strengths (Section 5.4).
@@ 5.1 Coverage of DQ dimensions by SHACL core
For each of the 18 DQ dimensions discussed in Zaveri et al. [1], we present SHACL core coverage as the normalized sum of coverage values assigned to each DQA metric: 1 for full coverage, 0.5 for partial coverage, and 0 for no coverage. The resulting sum is divided by the total number of metrics to obtain an overall coverage percentage. Figure 2 shows these results, where two dimensions with 100% coverage stand out: Representational Conciseness and Security. In these cases, the implemented metrics check for the presence or use of specific properties or classes.
Seven dimensions present a coverage between 50-90% : Timeliness, Understandability, Consistency, Completeness, Syntactic Validity, Interoperability, Interpretability, and Versatility.
Regarding Timeliness, SHACL core partially covers both of the proposed metrics. While it cannot directly compute the formulas these metrics rely on, like T1’s formula or T2’s time-distance calculation, it can still check related aspects. For example, for T1, we defined a shape that detects outdated entities via temporal annotations, and for T2, the shape verifies if the dataset is up to date, even though we cannot compute distances with SHACL core.
@@
5.2 Limitations of SHACL core
Despite its usefulness for validating RDF graphs, SHACL Core has limitations when applied to DQA:
@@ 5.2 Limitations of SHACL core
Node-Centric Scope. Constraints in SHACL core are evaluated for individual focus nodes and their immediate property values. This node-centric approach limits the ability to compute network-level measures, which are essential for evaluating the Interlinking dimension.
No Cross-Entity Comparison. SHACL core lacks mechanisms to compare property values across different entities in the graph. As a result, it cannot assess functional dependencies or detect duplicated entities, which are key for dimensions such as Semantic Accuracy and Conciseness.
No Arithmetic or Dynamic Expressions. SHACL Core lacks support for arithmetic operations and dynamic expressions. For instance, it cannot evaluate conditions like year = now() - birthDate. This limits the applicability of SHACL core for metrics that depend on computed values.
@@
Restricted Targeting Mechanism. SHACL provides only a small set of predefined target types (e.g., sh:targetClass, sh:targetNode). This makes it difficult to define metrics that need to assess patterns over arbitrary triples or general predicate-based constraints.

### v3

@@ Abstract
Knowledge graphs have been widely adopted in both enterprises, such as the Google Knowledge Graph, and open platforms like Wikidata, to represent domain knowledge and support artificial intelligence applications. They model real-world information as nodes and edges. To embrace flexibility, knowledge graphs often lack enforced schemas (i.e., ontologies), leading to potential data quality issues, such as semantically overlapping nodes. Yet ensuring their quality is essential, as issues in the data can affect applications relying on them. To assess the quality of knowledge graphs, existing works propose either high-level frameworks comprising various data quality dimensions without concrete implementations, define tools that measure data quality with ad-hoc SPARQL queries, or promote the usage of constraint languages, such as the Shapes Constraint Language (SHACL), to assess and improve the quality of the graph. Although the latter approaches claim to address data quality assessment, none of them comprehensively tries to cover all data quality dimensions. In this paper, we explore this gap by investigating the extent to which SHACL core can be used to assess data quality in knowledge graphs. Specifically, we defined SHACL shapes for 69 data quality metrics proposed by Zaveri et al. [1] and implemented a prototype that automatically instantiates these shapes and computes the corresponding data quality measures from their validation results. All resources are provided for repeatability.
@@
Knowledge graphs (KGs) have been increasingly used to represent domain knowledge and support artificial intelligence applications, such as information retrieval and question answering [2]. As a result, ensuring the quality of KGs is crucial for applications that rely on their input [3]. Major technology companies, such as Google (Google Knowledge Graph [4]) and Amazon (Alexa Knowledge Graph [5]), have developed proprietary KGs [2], while collaborative KGs, such as Wikidata [6], DBpedia [7] and YAGO [8], have emerged as open-source alternatives. This growing adoption of KGs is reflected in the growth of the Linked Open Data (LOD) cloud, which has enabled the publication of numerous datasets across domains, with 1 656 resources as of Nov. 2024 [9].
@@
In recent years, constraint languages, such as the Shapes Constraint Language (SHACL) [25] and Shapes Expression Language (ShEx) [26], have emerged to enable RDF graph validation by specifying constraints as shapes [10]. Despite syntactic differences, both languages enable the definition of constraints on nodes and their value nodes (i.e., values reachable via properties or paths), allowing for the detection of data violations [27]. These languages enable a low-level perspective on DQ, focusing on concrete error detection through constraint checks. Unlike ad-hoc SPARQL-based approaches, constraint languages like SHACL offer a formal way to express validation rules.
@@
This paper explores how SHACL core can be leveraged for DQA. In particular, we want to investigate the extent to which we can connect the high-level view on DQ dimensions with the low-level view on DQ, which entails identifying constraint violations using SHACL shapes. Thus, the contributions of this paper are as follows:
@@ 2 Related Work
In addition, several tools have been proposed over the years to assess the quality of KGs and linked data (which we consider a type of KG, though not all KGs follow Linked Data principles). Some of these tools focus on specific DQ dimensions [37, 38], while others try to cover a wider range of them [21, 23, 39, 40, 24, 41, 42, 43]. Moreover, some tools focus on assessing the quality of SPARQL endpoints [41, 38, 37], while others focus on the quality of the data itself [21, 23, 39, 40, 24, 42, 43]. Assessment is primarily done via ad-hoc SPARQL queries [23, 39, 37, 24, 41, 22, 43], while some introduce more complex techniques [21, 42]. Only [24] considers the usage of a constraint language. Additionally, some of these tools haven’t been maintained for some time or aren’t available. In Appendix A we present a table summarizing these tools.
@@ 3 SHACL shapes for DQ dimensions
This paper builds on the comprehensive DQ metrics survey by Zaveri et al. [1], which groups dimensions into four categories. We follow their structure, defining SHACL core shapes for each metric or explaining SHACL core’s limitations when shapes cannot be defined.111 A discussion on SHACL extensions and the justification for the focus on SHACL core can be found in Section 5.3. Across the paper we use SHACL interchangeably with SHACL core, unless stated otherwise. We present the results of this study in tables 1–4, which illustrate the feasibility of implementing metrics from [1] with SHACL core. Symbols , p, and x denote the degree to which metric implementation was possible (full, partial, or not at all, respectively). For the shape definition, we made some realistic assumptions (A1–3):
@@
The ontology used is sufficiently defined: each class c and property p has triples c,rdf:type,rdfs:Class and p,rdf:type,rdf:Property, respectively. Each property p, includes domain and range triples p,rdfs:domain,d and p,rdfs:range,r. Other property characteristics (e.g., irreflexivity, asymmetry) are also defined. For every instance i of a class c defined in the ontology, i,rdf:type,owl:NamedIndividual exists.
@@
We acknowledge that these assumptions may not always hold in real-world KGs. If (A1) is violated, untyped instances are excluded from validation. Likewise, when (A2) is not satisfied, schema characteristics such as range, domain, or property constraints (e.g., symmetry, irreflexivity) cannot be checked. One possible direction to relax these requirements is to mine such characteristics automatically or leverage profiling statistics to approximate range and domain information, thereby supporting the construction of an “initial” ontology. Regarding (A3), this is inherent to DQA, as some quality dimensions rely on domain knowledge, limiting their assessment. Tables 1–4, mark the use of assumptions with , where X is the assumption number. In the following subsections, we summarize SHACL core coverage per group, discuss a representative DQ metric for a single dimension of each group, and, if applicable, provide the corresponding shape in Turtle syntax. For the implemented shapes, we also indicate the DQ measure type: binary measure (1 if no violations, 0 otherwise), ratio measure (violation-based ratios), or composite measure (aggregated score across shape instances for specific properties or classes).
@@ 3.1 Accessibility
Performance refers to how efficiently a system that hosts a large dataset can process data. For this dimension, Zaveri et al. [1] defined four metrics. We showcase the assessment of Performance with metric P1 in SHLABEL:shape:performance_hash_uris_entities. P1 was defined by [44] and checks for slash-URIs in datasets with over 500 000 triples. According to W3C recommendations, slash-URIs are preferred to identify entities in large graphs because hash (“#”) URIs can cause performance issues [45]. We follow the W3C Best Practices [46] and apply this rule to entities’ URIs. Here, SHLABEL:shape:performance_hash_uris_entities targets subjects of triples with predicate rdf:type, applying a regex pattern that does not allow # to appear in URIs. Note that the regex pattern identifies any hash occurrence (not just at the end of the URI) to cover cases such as https://www.example.org#Example1/, which still loads all entities with https://www.example.org# as base namespace. Therefore, the validation result will output nodes whose URI contains a hash in any part of the URI, not just at the end.
@@ 4 Prototype
This section presents our prototype for DQA using SHACL core. It is built with the Python libraries rdflib and PySHACL and provides a Streamlit dashboard to visualize results. The code is available on GitHub at [47].
Overview of the DQ assessment process. Our DQA process takes as input the data graph to be evaluated, optionally a metadata file (either the VoID or DCAT description of the graph), and an ontology, along with a set of vocabularies used in the data graph. Without an ontology, vocabularies, or a metadata file, fewer shapes are instantiated, and some, such as those targeting void:Dataset or those checking for class/property labels, are not validated, as they rely on these artifacts. A configuration file allows users to customize preferred properties needed for instantiating SHACL shapes. We use, for example, rdf:type, rdfs:label, rdfs:comment, and owl:sameAs as default properties for defining shapes, but users may change this.
@@
After obtaining this information, we (2) instantiate the shapes’ templates and generate the shapes graph. Our prototype stores SHACL shapes as reusable templates with variables that are replaced with actual values during instantiation – a process where the template is populated with specific properties from either the graph profiling results or the configuration file of the dataset. For example, shape SHLABEL:shape:performance_hash_uris_entities uses the type property specified in the configuration file to create a concrete shape from its template form. Moreover, shape SHLABEL:shape:consistency_uniqueness_inverse_functional_property is instantiated with inverse-functional properties used in the dataset, obtained from the graph profiling results. Before shape validation, we perform a pre-processing step, based on assumption (A2), to enrich the data graph with necessary triples. Specifically, we add type declarations for OWL-defined properties and classes defined in ontology/vocabularies, marking each property as an rdf:Property and each class as an rdfs:Class. We also explicitly type instances in vocabularies/ontology as owl:NamedIndividual to distinguish data entities from schema elements. For terms defined as rdfs:Datatype, we type them as rdfs:Class, in line with the RDFS entailment rule rdfs11. Additionally, when multiple vocabularies are reused, we include all relevant axioms in the data graph, as stated by the SHACL specification [25]. Then, we use the validator (3) provided by the PySHACL library to validate the shapes against the data graph, ontology, vocabularies, or metadata file, depending on the shape. Once the validation is completed, we (4) calculate the DQ measures from the validation result, which are later stored in a CSV file. Finally, the results can be visualized in a dashboard (5).
@@
Shape instantiation details. For the 69 metrics defined in Zaveri et al. [1], we defined 64 shapes. The mapping is not one-to-one: some metrics have no shape, others (e.g., V2 in Section 3.4) have multiple. We identified 38 shapes that could be instantiated without domain expert input, and excluded five more: one requiring merged graphs (SH11) and 4 using non-standard vocabularies (S13, S14, S53, S55) (see Appendix B). Of the 38, 11 rely on vocabulary or ontology terms, and are instantiated only with those found in the data graph (e.g., SHLABEL:shape:consistency_uniqueness_inverse_functional_property uses only inverse-functional properties present in the data). This avoids generating unnecessary shapes, as many data graphs may partially reuse vocabularies. The exceptions are the shapes for metrics CN2 and CP1: CN2 checks for property/class misuse by instantiating with all available classes and properties since misused ones do not appear in profiling results, while CP1 uses all defined classes in the vocabulary to check if they are used. See Appendix C for additional aspects of shape instantiation aimed at improving runtime efficiency.
@@
Evaluation. We evaluated the prototype with three datasets from the LOD Cloud222All datasets are available at https://zenodo.org/records/16644385: Temples of the Classical World (15 326 triples and 1 363 entities), DBTunes - John Peel Sessions (271 369 triples and 76 056 entities), and Drugbank (3 646 181 triples and 316 555 entities). Validation time ranged from 40 seconds to 3 hours for 500–740 instantiated shapes, depending on the dataset. Longer times are likely due to both more triples and more shapes. As the prototype aimed to test the SHACL-based DQA approach, performance aspects were left outside its scope, and no runtime experiments were conducted since validation relied on an external library. Table 5 presents the assessment results for the dataset Temples of the Classical World in detail. The results for the other datasets are provided on GitHub at [47]. While Temples of the Classical World demonstrates strengths in aspects such as URI design and labeling, it exhibits several quality issues, particularly in metadata aspects, and correct vocabulary usage.
@@ 5 Suitability of SHACL core to assess RDF data quality
We now discuss the suitability of SHACL core to assess RDF DQ. Section 5.1 highlights where SHACL core falls short in covering certain DQ dimensions – metrics not mentioned are considered covered (see Section 3 and Appendix B). We then examine SHACL core’s strengths and limitations (Section 5.2), and discuss SHACL core’s extensions (Section 5.3).
@@ 5.1 Coverage of DQ dimensions by SHACL core
For each of the 18 DQ dimensions discussed in Zaveri et al. [1], we present SHACL core coverage as the normalized sum of coverage values assigned to each DQ metric: 1 for full coverage, 0.5 for partial coverage, and 0 for no coverage. The resulting sum is divided by the total number of metrics to obtain an overall coverage percentage. Figure 2 shows these results, where two dimensions with 100% coverage stand out: Representational Conciseness and Security. In these cases, the implemented metrics check for the presence or use of specific properties or classes.
Seven dimensions present a coverage between 50-90% : Timeliness, Understandability, Consistency, Completeness, Syntactic Validity, Interoperability, Interpretability, and Versatility. Regarding Timeliness, SHACL core partially covers both of the proposed metrics. While it cannot directly compute the formulas these metrics rely on, like T1’s formula or T2’s time-distance calculation, it can still check related aspects. For example, for T1, we defined a shape that detects outdated entities via temporal annotations, and for T2, the shape verifies if the dataset is up to date, even though we cannot compute distances with SHACL core.
@@
5.2 Strengths and Limitations of SHACL core
SHACL core has many strengths for specific aspects of DQA. For example, it can perform syntactic validation, including pattern validation, datatype checks, and verifying whether values fall within specified ranges or lists. It can validate consistency between the data and vocabulary or ontology definitions, such as ensuring correct use of a property’s domain or range. It can also support best practices for publishing Linked Data, such as the use of hash URIs for entities, labels in resources (i.e., entities, classes, and properties), and some characteristics of URI designs (i.e., short URIs and no parameters). Moreover, SHACL can verify the correct application of property characteristics such as irreflexive, functional, asymmetric, and inverse-functional. Finally, it is well-suited for defining the expected structure of class instances, specifying required properties, their types, and other basic constraints [13].
However, its applicability to DQA is still limited in several ways:
@@ 5.2 Strengths and Limitations of SHACL core
Node-Centric Scope. Constraints in SHACL core are evaluated for individual focus nodes and their immediate property values. This node-centric approach limits the ability to compute network-level measures, which are essential for evaluating the Interlinking dimension (covered by SHACL-SPARQL).
No Cross-Entity Comparison. SHACL core lacks mechanisms to compare property values across different entities in the graph. As a result, it cannot assess functional dependencies or detect duplicated entities, limiting its applicability for Semantic Accuracy and Conciseness (covered by SHACL-SPARQL).
No Arithmetic or Dynamic Expressions. SHACL core lacks support for arithmetic operations and dynamic expressions. For instance, it cannot evaluate conditions like age = now() - birthDate. This limits the applicability of SHACL core for metrics that depend on computed values, such as Timeliness and Trustworthiness (covered by SHACL-SPARQL).
@@
Restricted Targeting Mechanism. SHACL provides only a small set of predefined target types (e.g., sh:targetClass, sh:targetNode). This makes it difficult to define metrics that need to assess patterns over arbitrary triples or general predicate-based constraints (covered by SHACL-SPARQL).